Suppose cost of production of a firm is given by:





What amount of the product should be produced to maximize profit of the firm if price per unit output is 74? Find the maximum profit or loss.



What quantity of the product should be produced if price per unit is 40?



Find the maximum profit. What will be the decision of the firm to stay in production or stop operation? Why?



What will be the decision of the firm if price per unit is 10?



What is the shut down price level?



Determine the supply function of the firm with domain or range of out put

1
Expert's answer
2022-02-14T14:46:51-0500

Let

"TC=Q^2+2Q+1"

P=MC


P=74

MC=2Q+2

74=2Q+2

72=2Q

Q=36

"Profit=TR-TC=36\times74-(36^2+2\times36+1)=2664-1369=1295"


P=40

MC=2Q+2

40=2Q+2

38=2Q

Q=19

"Profit=TR-TC=40\times19-(19^2+2\times19+1)=760-400=360"

TR>TC the company continues to operate


P=10

10=2Q+2

Q=4

"Profit=TR-TC=4\times10-(4^2+2\times4+1)=40-25=15"


MC>MR

Q=0


(0;36) and above 36